OpenStudy (anonymous):

how can we find the impulse response of a moving average system in discrete time & how can we find the output of a given system if we know its input as impulse????

OpenStudy (anonymous):

Translate the expression into a word phrase. k + 12 A. a number less than 12 B. 12 more than a number C. the product of a number and 12 D. the quotient of a number and 12

OpenStudy (anonymous):

Signals And Systems?? Convolution?

OpenStudy (anonymous):

DSP not S&S but yup its related 2 convolution

OpenStudy (anonymous):

excuse me ... help me.. m in great trouble...

OpenStudy (anonymous):

\[y[n]=1/5 \sum_{4}^{0} x(n-k)\]

OpenStudy (anonymous):

is the relationship b/w input & output n we have to find impulse response of MA system

OpenStudy (anonymous):

MA??

OpenStudy (anonymous):

moving average system

OpenStudy (anonymous):

one million

OpenStudy (anonymous):

??? means

OpenStudy (anonymous):

MA = one million

OpenStudy (anonymous):

\[x[n]=\cos(w0n)\]

OpenStudy (anonymous):

didn't get ur rep... sumanta

OpenStudy (anonymous):

\[y[n]=\frac{1}{5} \sum_{k=4}^{0} x(n-k)\] Right??

OpenStudy (anonymous):

how it can be 4 to 0?? -4 to 0??

OpenStudy (anonymous):

Write your question properly.. :)

OpenStudy (anonymous):

sorry its 0 to -4

OpenStudy (anonymous):

:S

OpenStudy (anonymous):

answer plz...

OpenStudy (anonymous):

\[y[n]=\frac{1}{5} \sum_{k=-4}^{0} x(n-k)\]

OpenStudy (anonymous):

For Impulse Response: Replace input \(x[n]\) by \(\delta[n]\), and \(y[n]\) by \(h[n]\)..

OpenStudy (anonymous):

\[h[n]=\frac{1}{5} \sum_{k=-4}^{0} \delta(n-k)\]

OpenStudy (anonymous):

just this is the answer ???

OpenStudy (anonymous):

NO..

OpenStudy (anonymous):

Evaluate this.. :P And I don't know how to do that..

OpenStudy (anonymous):

n again sorry.. m in great stress.. limits are 0 to +4...

OpenStudy (anonymous):

u dnt know??

OpenStudy (anonymous):

\[h[n]=\frac{1}{5} \sum_{k=0}^{4} \delta(n-k)\]

OpenStudy (anonymous):

Here, I know one formula, but I don't know to apply that: \[\sum_{k=0}^{\infty} \delta(n-k) = u[n]\]

OpenStudy (anonymous):

Moreover, it can be: \[h[n] = \frac{1}{5}(\delta(n) + \delta(n-1) + \delta(n-2) + \delta(n-3) + \delta(n-4))\]

OpenStudy (anonymous):

\[h[n] = \{\underline{\frac{1}{5}}, \frac{1}{5}, \frac{1}{5}, \frac{1}{5}, \frac{1}{5}\}\]

OpenStudy (anonymous):

\(n=0\) at first \(\frac{1}{5}\), then n = 1, 2,3, 4..

OpenStudy (anonymous):

I am not sure, if I am true, just I have solved it as my Heart guided me.. :P

OpenStudy (anonymous):

i think ur brain guided u ri8.. ys its the correct ans.. i have cnfmed it ..

OpenStudy (anonymous):

No, actually it was my Heart not brain, that's why I took a hell of time to solve it, if it would have been a brain, then it would have been done in a jiffy.. :P

OpenStudy (anonymous):

Actually, Digital Signal Processing is a subject which bases itself upon Signals And Systems, so as I am doing Signals And Systems, so may be I can help you.. :) But you have to give me time, I am no expert in it, I will take my time to tell you.. :) Because, I don't want to tell anyone here a wrong thing ever.. :)

OpenStudy (anonymous):

I have posted this post's link on the post where you asked about Time Period of Discrete time, so please delete that, I can't do that because my net is taking revenge from me.. :) It is going slow like a turtle.. :)